# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

### v4

@@ C.3 Different Choice of $\lambda$ in UCB
Figure 13: Runtime optimization performance with different exploration level $\lambda$ in UCB.
@@
We examined how the exploration parameter $\lambda$ in UCB influences the BO performance in high-dimensional optimization. To this end, we used the Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}. We conducted MLE training via ADAM optimization. We show the ultimate and runtime optimization performance in Fig. 13, respectively. It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close. However, on DNA, SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$. This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
C.4 Alternative Acquisition Functions
@@
where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data. We used BOTorch to maximize EI through L-BFGS. The most recent work [1] introduced log-EI to overcome the numerical challenges in optimizing the original EI acquisition function. Accordingly, we also tested log-EI using the BoTorch implementation. To conduct TS at each step, we used Sobol Sequence [28] to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior. We then selected the input with the largest function value as the next query. We denote this method by TS-3K. To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence. We denote this method as TS-10K. We used BOTorch implementation of both methods. We used the Matérn kernel and MLE training at each step.
@@
The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig. 14 and 15, respectively. As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline. On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst. The failure of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work [1]. The failure of the TS might be due to the high-dimension of the input space. A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function. The use of log-EI performed excellently, achieving results comparable to those obtained with UCB.
Appendix D Limitation
@@
Our current work has identified gradient vanishing as a critical failure mode of SBO in high-dimensional settings. We have proposed an effective, robust initialization method to mitigate this issue. However, we have not yet analyzed whether there is an upper limit to the input dimension for which SBO can function properly, or what that limit might be. To answer these questions, we will continue our investigation in future work, both theoretically and empirically.

### v5

@@ C.2 Different Choice of $\lambda$ in UCB
Figure 8: Runtime optimization performance with different exploration level $\lambda$ in UCB. In the main paper and Figure 5 and 6, SBO was compared with other methods using $\lambda=1.5$.
@@
We examined how the exploration parameter $\lambda$ in UCB influences the BO performance in high-dimensional optimization. To this end, we used the Maérn kernel and varied $\lambda$ from {0.5, 1.0, 1.5, 2.0}. We show the runtime optimization performance in Figure 8. It can be seen that, in most benchmarks (Mopta08, Rover, NAS201, DNA, etc.), the performance with different $\lambda$’s is close. However, on SVM and Ackley, the performance of $\lambda=0.5$ is typically worse than the bigger choice of $\lambda$. This might be due to that those problems are more challenging (e.g., Ackley has many local minima), a larger exploration level ($\lambda$) is needed to identify better optimization outcomes.
C.3 Alternative Acquisition Functions
@@
where $\Psi(\cdot)$ and $\Phi(\cdot)$ are the CDF and PDF of the standard Gaussian distribution, ${\bf x}^{*}$ is the best sample that gives the largest function value among all the queries up to the current step, $\mu({\bf x})$ and $\sqrt{v({\bf x})}$ are the posterior mean and standard deviation of $f({\bf x})$ given current training data. We used BOTorch to maximize EI through L-BFGS. The most recent work (Ament et al., 2024) introduced log-EI to overcome the numerical challenges in optimizing the original EI acquisition function. Accordingly, we also tested log-EI using the BoTorch implementation. To conduct TS at each step, we used Sobol Sequence (Sobol’, 1967) to sample 3K inputs in the domain, and jointly sampled the 3K function values from the GP posterior. We then selected the input with the largest function value as the next query. We denote this method by TS-3K. To explore more input candidates, we also employed the Lancos and conjugate gradient descent to approximate the posterior covariance matrix, and sampled 10K input candidates from Sobol Sequence. We denote this method as TS-10K. We used BOTorch implementation of both methods. We used the Matérn kernel.
@@
The ultimate and runtime optimization performance for using EI, TS-3K and TS-10K, are shown in Fig. 9 and 10, respectively. As we can see, in most cases, such as Mopta08 (128), SVM (388), DNA (180) and Rosenbrock (100, 100), both EI and TS perform worse than UCB, and also worse than the best baseline. On Ackley(150, 150) and Ackley(300, 150), the performance of EI and TS are even among the worst. The failure of EI might arise from the gradient vanishing during the optimization, as pointed by the recent work (Ament et al., 2024). The failure of the TS might be due to the high-dimension of the input space. A few thousand candidates might not be sufficient to cover the whole space; as a result, the selected queries are inferior in indicating the best direction to optimize the target function. The use of log-EI performed excellently, achieving results comparable to those obtained with UCB.
Appendix D Length-scale Initialization Strategies Across Different Gaussian Process Libraries
In this section, we have investigated the initialization strategy of a (non-exhaustive) list of GP training libraries. Most libraries employ a small constant initialization while a few others created the initialization based on the distances between the training data points.
D.1 Constant Initialization
D.2 Data-Dependent Initialization
  • DiceKriging [^10] https://github.com/cran/DiceKriging initializes the length-scale randomly based on the range of the input features. The initial length scale for each feature is sampled from a uniform distribution, with a lower bound of $1\times 10^{-10}$ and an upper bound set to twice the maximum difference of the variable values in the training dataset.
  • hetGP [^11] https://github.com/cran/hetGP initializes the length-scale to the 50th percentile of the pairwise distances between the training data points.
  • GPVecchia [^12] https://github.com/katzfuss-group/GPvecchia initializes the length-scale to be proportional to the mean of the pairwise distances between the training data points.
Appendix E Limitation
@@
Our current work has identified gradient vanishing as a critical failure mode of SBO in high-dimensional settings. We have proposed an effective, robust initialization method to mitigate this issue. However, we have not yet analyzed whether there is an upper limit to the input dimension for which SBO can function properly, or what that limit might be. To answer these questions, we will continue our investigation in future work, both theoretically and empirically. In addition, we empirically found that the choice of optimizer can also impact GP training performance in high-dimensional settings. For instance, L-BFGS tends to be less stable than ADAM, while RMSProp can outperform ADAM in certain cases. In future work, we will continue exploring initialization strategies to enhance the robustness of GP training across different optimizers.